OpenStudy (anonymous):

Help! Fan + Medal Consider the function y = 3x5 – 25x3 + 60x + 1. Find the critical points of this function. None –2, –1, 0, 1, 2 –2, –1, 1, 2 1, 2 –2, 2

OpenStudy (p0sitr0n):

you find the critical points of a function (extremums) by taking f'(x)=0

OpenStudy (anonymous):

I think its B but I want to make sure, you know

OpenStudy (anonymous):

But I tend to suck at taking derivatives of functions thats why I always want someone else's opinion

OpenStudy (p0sitr0n):

f'(x)=15x^4-75x^2+60=0 then solution space is : x =+-1, +-2

OpenStudy (p0sitr0n):

use the power rule \[\frac{ \delta x^n }{ \delta x } = n x^{n-1}\]

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

So I got -2,-1,1,2 again so I am guessing that's the right answer. Thanks :]